OpenStudy (anonymous):

solve the equation -3(x-5)=-5(-3-4x)

OpenStudy (anonymous):

help plzzz

OpenStudy (anonymous):

-3x = 20x or x=0 im not exactly sure

OpenStudy (anonymous):

thank you i got it

OpenStudy (kamille):

Well, firstly you need to -3(x-5)=-5(-3-4x) -3x+15=15+20x -3x-20x=0 x=0

OpenStudy (anonymous):

thank you guys